New vidyo room:

News:

IRIS database per il gruppo di ricerca da riempire. Selezionare le pubblicazioni in cui abbiamo contribuito. Da inserire nel google form di Paolo.

Framework & production

no news are good news.

Analysis

B physics

Alberto: writing thesis. Martino nominated for the HFLAV Enrico: article reviewers asked about the correlations of the syst uncertainties. We have a method.

Heavy composite neutrinos

MAtteo: writing the thesis. YEsterday pre-approval meeting. Conditionally preapproved, minor fixing, nothing dramatic. Good feedback from the ARC members (Greg and Emanuela Barberis). We can unblind the first 4 bins of the signal region and see how it goes.

VBS

no news.

WH_SS

Alberto: meeting last week. writing his thesis. Analysis was preapproved.

Same sign dileptons

no news

Htomumu/Htocharm

Pier: just finished working on the analysis. Tail of the analysis on the muons. working on the refit of the beam spot. Pier: work on the merged topology (2 charms making a fat jet). targeting full run2 for Moriond next year.

EPR

please pledge.

Detector

Trigger

Jacopo: interacting w/ Antonio and Marco B. Attached one OBDT to a chamber. Taken data. Seen features (duplicated hits) to understand.

Timing Layer

the assegno tecnologico is open. we have 2 candidated. The radiation tests in Legnaro for now moved to december. Paolo Meridiani contacted us to make tests on Peltier cooling sipm. RobertoC: the dark current of the Sipm can not work at -30C. And you want to anneal them in idle periods. This can be done with Peltier. The main group is from Vanderbilt at CERN. Both tests and simulations. Workshop on the peltier, Friday on nov 20th. funds need to be found (0.5-1MCHF). Supported by referees.

DT

Advancing on the studies on the aging. We can separate 2 origin of the aging. 1 glue. 2 mylar. The second component is not always present. Seems that the rate of gas change (not gas flow) is affecting the aging. MB2 wire extracted (with small charge, about 1/100 of HLLHC) had no deposit on it.

Tracker/tracking

Pier: in contact with Nicola, planning the testing of the chip. For now kind of idle, waiting for chips from FNAL. Mia: run3 enrico and otr ppl working on ghte miniAOD format.

A.O.B.

notte della ricerca. -- RobertoRossin - 2020-11-10
Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2020-11-10 - RobertoRossin
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback